If you deposit $1800 in an account paying 2.3% interest for "t" years, what is the equation for the amount of money at time "t"? When will the account have $2500?
Answer by fractalier(6550)

If this is simple interest, then
interest I = Prt
The total will be P+I at that point, t years in the future...so
Total = 1800 + 1800(.023)t
Now if we set that equal to 2500, we get
2500 = 1800 + 1800(.023)t
700 = 41.4t
t = 700 / 41.4 = 16.9 years